OpenStudy (dtan5457):

Suppose h(2z+3)=(5-2z)/(4+z) find h(-9) find h^-1(4) find h(z)

OpenStudy (dtan5457):

Anyone know what type of math this called? I never seen this

OpenStudy (dtan5457):

@DanJS @dan815 @Loser66

OpenStudy (firekat97):

do you by any chance have answers to these?

OpenStudy (dtan5457):

no

OpenStudy (loser66):

I don't know whether it works or not, just give it a try; Let t = 2z +3, then \(z = \dfrac{t-3}{2}\) So \(h(2z+3) = h (t) =\dfrac{5-(t-3)}{4+\dfrac{t-3}{2}}= \dfrac{16-2t}{5-t} \)

OpenStudy (loser66):

So as h(z) , that is \(h(z) = \dfrac{16-2z}{5-z}\)

OpenStudy (loser66):

From it, we can find h(-9), \(h^{-1} (4)\) is just let it =4, solve for z

OpenStudy (anonymous):

If you're looking for more information on this type of problem, it's called a functional equation. It may or may not be a part of functional analysis, I'm not sure.

OpenStudy (dtan5457):

its alright, i figured these 3 out.